### v1

# Is Parameter Collision Hindering Continual Learning in LLMs?

Abstract
Large Language Models (LLMs) often suffer from catastrophic forgetting when learning multiple tasks sequentially, making continual learning (CL) essential for their dynamic deployment. Existing state-of-the-art (SOTA) methods, such as O-LoRA, typically focus on constructing orthogonality tasks to decouple parameter interdependence from various domains. In this paper, we reveal that building non-collision parameters is a more critical factor in addressing CL challenges. Our theoretical and experimental analyses demonstrate that non-collision parameters can provide better task orthogonality, which is a sufficient but unnecessary condition. Furthermore, knowledge from multiple domains will be preserved in non-collision parameter subspaces, making it more difficult to forget previously seen data. Leveraging this insight, we propose Non-collision Low-Rank Adaptation (N-LoRA), a simple yet effective approach leveraging low collision rates to enhance CL in LLMs. Experimental results on multiple CL benchmarks indicate that N-LoRA achieves superior performance ($\bm{+2.9\%}$), higher task orthogonality (×$\bm{\times 4.1}$ times), and lower parameter collision (×$\bm{\times 58.1}$ times) than SOTA methods.
Is Parameter Collision Hindering Continual Learning in LLMs?
Shuo Yang1∗, Kun-Peng Ning1∗, Yu-Yang Liu1†, Jia-Yu Yao1, Yong-Hong Tian1, Yi-Bing Song2, Li Yuan1† 1School of Electronic and Computer Engineering, Peking University 2DAMO Academy, Alibaba Group {shuo_yang, ningkp}@stu.pku.edu.cn, {liuyuyang13, jiayu_yao, yhtian, yuanli-ece}@pku.edu.cn yibingsong.cv@gmail.com
These authors contributed equally to this work.Corresponding authors.
1 Introduction
Continual Learning (CL) De Lange et al. (2021); Silver et al. (2013) is essential for adapting large language models (LLMs) to new tasks, but it often leads to catastrophic forgetting McCloskey and Cohen (1989), where knowledge from previous tasks is overwritten. Various methods, such as rehearsal-based Rebuffi et al. (2017), regularization-based Kirkpatrick et al. (2017), and architecture-based approaches Mallya and Lazebnik (2018), have been proposed to mitigate this, but they significantly increase computational demands due to the vast number of parameters in LLMs  Wu et al. (2024). Recent advances in parameter-efficient fine-tuning (PEFT), such as LoRA Hu et al. (2021), introduce a minimal set of trainable parameters while maintaining the rest of the pre-trained LLM fixed. However, the task-specific parameters may collide with each other, potentially overwriting earlier knowledge and thus leading to poor model performance.
Figure 1: (a) Orthogonal but Parameter Collision: Tasks $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$ are mutually orthogonal but interaction within each space, resulting in parameter collision. (b) Non-collision and Orthogonal: Tasks $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$ update only along distinct, non-conflicting subspaces, preserving prior task knowledge. (c) Performance Comparison: N-LoRA (red) and O-LoRA (blue) are compared across various metrics, with N-LoRA achieving lower collision rates, improved orthogonality, and superior average accuracy.
To tackle this challenge in LLMs, existing studies Farajtabar et al. (2020); Saha et al. (2021); Wang et al. (2023a) attempted to employ orthogonal gradient descent-based methods to decouple parameter interdependence across different domains. Among them, O-LoRA (current SOTA CL method in LLMs) incrementally updates task-specific parameters by constraining them to the orthogonal subspace of previous tasks, ensuring that the LoRA parameters learned from past tasks remain fixed and independent across different subspaces Wang et al. (2023a). However, despite this independence, LoRA parameters within different subspaces may still interact, leading to potential parameter collisions.
In this paper, we reveal that constructing non-collision LoRA parameters provides a more effective solution to the problem of parameter interdependence, as non-collision serves as a sufficient but not necessary condition for orthogonality. As illustrated in Figure 1(a), assuming a three-dimensional parameter space, existing methods (e.g. O-LoRA) aim to learn three mutually orthogonal (blue box) task vectors $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$. Each task vector $[x_{i},y_{i},z_{i}]$ occupies all three subspaces of the parameters, and these parameters will interact with each other (red box), hurting the model performance. In contrast, as illustrated in Figure 1(b), non-collision constructs an extremely sparse parameter space, allowing each task vector to be updated in a local subspace without affecting each other. Non-collision parameters naturally achieve better task orthogonality (blue box), effectively decoupling parameter interdependence while preserving prior knowledge.
Leveraging the above insights, we propose Non-collision Low-Rank Adaptation (N-LoRA), a simple yet effective approach to tackling the continual learning problems in LLMs. Instead of directly optimizing the collision rate, we introduce $\ell_{1}$ constraints on task-specific LoRA parameters to make them extremely sparse, thereby significantly reducing the collision and dependency between various task vectors. As illustrated in Figure 1(c), compared with the orthogonal-based method (O-LoRA), the proposed non-collision-based approach (N-LoRA) can significantly achieve better task orthogonality (×$\bm{\times 4.1}$ times), lower parameter collision (×$\bm{\times 58.1}$ times), and superior average accuracy ($\bm{+2.9\%}$) in all cases.
We empirically validate the effectiveness of the proposed N-LoRA through extensive experiments, demonstrating that it significantly outperforms prior SOTA methods across various datasets and model architectures. The contribution of this paper can be summarized as follows:
  • We reveal the intrinsic relationship between parameter orthogonality and collisions, clarifying that the key to addressing catastrophic forgetting lies in preventing parameter collisions, not just enforcing orthogonality.
  • We introduce a simple yet effective method called N-LoRA for tackling the CL problem in LLMs, which leverages low collision parameters to enhance the model performance.
  • Our experiments on various CL benchmarks demonstrate that N-LoRA not only surpasses previous SOTA methods but also achieves better orthogonality and a lower collision rate.
  • N-LoRA is compatible with existing PEFT-based approaches and serves as a plug-and-play solution that can substantially enhance the performance of existing SOTA methods.
2 Related Works
Parameter Efficient Fine-Tuning. Parameter Efficient Fine-Tuning (PEFT) He et al. (2021); Ding et al. (2022) enhances large language models (LLMs) performance by adjusting only a small portion of parameters, achieving results comparable to full fine-tuning while reducing computational costs. Key methods include Adapter tuning Houlsby et al. (2019), Prompt-based tuning Wei et al. (2021), P-tuning Liu et al. (2021a), Prefix tuning Li and Liang (2021), BitFit Zaken et al. (2021), and LoRA Wang et al. (2023a). In this work, we introduce a simple yet effective method built on LoRA to address this challenge, demonstrating strong performance.
Conventional Continual Learning. Conventional CL typically falls into three categories: (i) Rehearsal-based approaches Li and Hoiem (2017); Liu et al. (2021b); Lopez-Paz and Ranzato (2017) store and retrain on previous task data but may pose privacy concerns. (ii) Regularization-based approaches Kirkpatrick et al. (2017); Zenke et al. (2017) mitigate forgetting by imposing constraints that discourage significant changes in key weights. (iii) Architecture-based approaches Mallya and Lazebnik (2018); Wang et al. (2023b) expand model capacity dynamically or isolate existing weights with task-specific parameters to minimize interference between new and old tasks.
PEFT-Based Continual Learning for LLMs. LLMs with their vast number of parameters, create a significant computational burden for continual learning. To address this , methods like Learning to Prompt (L2P) Wang et al. (2022b) use a pool of small, learnable prompts to guide pre-trained models through tasks, while DualPrompt Wang et al. (2022a) combines expert and general prompts for task-specific and task-agnostic instructions. PEGP Qiao et al. (2024) utilizes Singular Value Decomposition(SVD) of task data to control gradient updates through orthogonal gradient projection, ensuring that new parameters do not interfere with the established feature space. O-LoRA Wang et al. (2023a) updates task-specific parameters by constraining them to an orthogonal subspace while keeping previous task parameters fixed However, O-LoRA struggles with parameter collisions, limiting its effectiveness in continual learning. Our proposed N-LoRA not only achieves better orthogonality than O-LoRA but also reduces parameter collisions and delivers superior performance.
3 Methodology
3.1 Continual Learning for LLMs
In this paper, We focus on continual learning for LLMs, where the model adapts sequentially to new tasks presented in a streaming manner without access to previous task data.
Orthogonality-Based CL for LLMs. Continual learning in LLMs poses computational challenges due to their large parameter space. To mitigate this, existing studies have used PEFT methods and employed orthogonal gradient descent-based approaches to decouple parameter interdependence across different tasks. Among these, O-LoRA introduces matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ based on the low-rank assumption Wang et al. (2023a), updating the parameters as $W=W_{0}+\Delta W=W_{0}+AB$. O-LoRA enforces orthogonality between the task subspaces by imposing constraints on $A_{t}$ and $A_{i}$, introducing an additional orthogonality loss:
$L_{\text{orth}}\left(A_{t}\right)=\sum_{i=1}^{t-1}\left\|A_{i}^{T}A_{t}\right\|^{2}$ (1)
However, despite the enforced orthogonality constraints, the parameter subspaces used by different tasks may still overlap, preventing the achievement of full orthogonality (orthogonality loss fails to converge). Additionally, LoRA parameters within different subspaces may interact, leading to potential parameter collisions, which in turn degrade the performance across tasks.
3.2 Orthogonality vs. Non-Collision
We reveal that non-collision LoRA parameters provide a more effective solution to parameter interdependence and show that non-collision is a sufficient but not necessary condition for orthogonality (see Appendix 7.1 for proof). The detailed analysis is as follows,
Definition 1.
For two parameter matrices $\Delta W_{1}$ and $\Delta W_{2}$ of the same dimensions, they are defined as non-collision if for every position $(a,b)$, it holds that $\Delta W_{1}[a,b]=0$ or $\Delta W_{2}[a,b]=0$.
Definition 2.
For two parameter matrices $\Delta W_{1}$ and $\Delta W_{2}$ of the same dimensions, they are said to be orthogonal if their matrix product satisfies $\Delta W_{1}^{\top}\Delta W_{2}=0$, where $\Delta W_{1}^{\top}$ denotes the transpose of $\Delta W_{1}$.
Theorem 1.
For two parameter matrices $\Delta W_{1}$ and $\Delta W_{2}$ of the same dimensions, non-collision is a sufficient but not necessary condition for orthogonality. Specifically,
$\forall(a,b),\quad\Delta W_{1}[a,b]=0\vee\Delta W_{2}[a,b]=0$
$\implies\Delta W_{1}^{\top}\Delta W_{2}=0,$
Obviously, non-collision parameters inherently result in orthogonality. In other words, reducing parameter collisions is a more efficient way to achieve better task orthogonality.
[…]
Visualization on Parameter Collisions. Based on the above analysis, we aim to visualize the relationship between parameter orthogonality and non-collision. We conducted experiments on the T5-large model, comparing parameter distributions using O-LoRA (orthogonality-based) and N-LoRA (non-collision-based) in a continual learning setup for Task 1 and Task 2. In the visualizations, parameters near zero are shown in gray, non-zero parameters in green and blue, and collisions in red.
[…]
3.3 Non-collision Low-Rank Adaptation
Building on the motivation above, the core goal of tackling continual learning in LLMs is to minimize parameter collisions as much as possible, i.e.,
$\min\sum_{\forall(a,b)}\mathbf{1}\{\Delta W_{1}[a,b]\neq 0\land...\land\Delta W_{n}[a,b]\neq 0\},$
where $\mathbf{1}\{\cdot\}$ is the indicator function. To achieve this, we propose to sparsify each LoRA parameter $\Delta W_{i}$, i.e.,
$\max\sum_{\forall(a,b)}\mathbf{1}\{W_{i}[a,b]=0\}$ (2)
We attempt to sparsify individual LoRA parameters to minimize overall collisions. We show that the collision rate will decrease quadratically with the sparsity rate in Appendix 7.2. To further simplify, we introduce a sparsity constraint directly on $\Delta W_{i}$ while freezing the parameters from previous tasks. Specifically, for a sequence of tasks $T=\{t_{1},t_{2},\dots,t_{n}\}$, a corresponding sequence of task-specific matrices $\Delta W=\{\Delta W_{1},\Delta W_{2},\dots,\Delta W_{n}\}$ is obtained through continual fine-tuning using LoRA. For the $i$-th task, the matrix $\Delta W_{i}=A_{i}B_{i}$ is subject to $\ell_{1}$ regularization:
$L_{\text{sparse}}=\lambda\ell_{1}(\Delta W_{i})=\lambda\lVert\Delta W_{i}\rVert_{1}$ (3)
where $\lambda$ is the sparsity hyperparameter. For each new task, only the task-specific LoRA parameters are updated, while the pre-trained model and previously learned LoRA parameters remain fixed. Consequently, the overall loss function for the $i$-th task is defined as:
$L=L_{\text{task}}+L_{\text{sparse}}=L_{t_{i}}+\lambda\lVert\Delta W_{i}\rVert_{1}$ (4)
Additionally, since the parameters from previous tasks are not needed during training or inference, they can be merged with the pre-trained model’s parameters, effectively reducing memory usage and maintaining the model’s efficiency.
Figure 3: Nuclear norms of O-LoRA and N-LoRA. The top plot (blue) shows the nuclear norms for O-LoRA, indicating the subspace dimensionality used by each layer. The bottom plot (red) shows the nuclear norms for N-LoRA. The dashed lines in both plots represent the average nuclear norm across all layers.
Discussion: N-LoRA can effectively reduce the dimension of task subspaces. Beyond reducing collision rates, sparsity offers the additional advantage of effectively reducing the dimensionality of task subspaces. The nuclear norm, which provides an upper bound on the rank of a matrix Li and Chen (2021), corresponds to the number of subspaces utilized by each adapter in PEFT-based continual learning. Taking the first task of the Standard CL Benchmark as an example, we computed the nuclear norm of the LoRA weights inserted at each layer of the T5-large model for both O-LoRA and N-LoRA, as illustrated in Figure 3. With the rank set to 8 (i.e., 8 subspaces per task), O-LoRA (top plot) nearly fully occupied these subspaces, averaging 6.80 across all layers. In contrast, N-LoRA, by leveraging sparsity, utilized an average of only 0.14 subspaces. This clearly demonstrates that N-LoRA significantly reduces the dimensionality of task subspaces. Furthermore, knowledge from multiple domains will be preserved in these extremely low-dimensional (non-collision) parameter subspaces, making it more difficult to forget previously seen data.
4 Experiments
4.1 Experimental Setups
Datasets. To evaluate our method, we use the Standard CL Benchmark Zhang et al. (2015) with five text classification datasets and the Large Number of Tasks Benchmark Razdaibiedina et al. (2023) encompassing 15 datasets, including GLUE Wang (2018), SuperGLUE Wang et al. (2019), and IMDB Maas et al. (2011). We follow the experimental setups of LFPT5 Qin and Joty (2021) and O-LoRA Wang et al. (2023a), using three different orders for each benchmark. Detailed task information and configurations are in Appendix 7.3, 7.4, and 7.5.
Metrics. Let $a_{i,j}$ be the test accuracy of the $i$-th task after training on the $j$-th task. $A_{i}$ denotes the LoRA $A$ matrix, and $\Delta W_{i}=A_{i}B_{i}$ matrix, both after training on the $i$-th task. We evaluate the model using the following six metrics:
(\@slowromancapi@) Average Accuracy (AA) Lopez-Paz and Ranzato (2017): The average accuracy of all tasks after training on the final task, defined as:
$A_{T}=\frac{1}{T}\sum_{i=1}^{T}a_{i,T}$ (5)
(\@slowromancapii@) Forgetting Rate (F.Ra) Chaudhry et al. (2018): This metric evaluates the amount of knowledge forgotten across the first $T-1$ tasks, as:
$F_{T}=\frac{1}{T-1}\sum_{t=1}^{T-1}\left(\max_{k=t}^{T-1}a_{k,t}-a_{T,t}\right)$ (6)
(\@slowromancapiii@) Orthogonal Overlap (OO) Wang et al. (2023a): Follow the O-LoRA, it quantifies the orthogonal overlap between the subspace of the $T$-th task and the subspaces of the previous $T-1$ tasks, calculated as:
$O_{T}=\sum_{i=1}^{t-1}\left\|A_{i}^{T}A_{t}\right\|^{2}$ (7)
(\@slowromancapiv@) Adapter Weight Orthogonality Magnitude (AWOM) Bhardwaj et al. (2024): This metric measures the degree of orthogonality between adapter weights, defined as:
$AWOM_{T}=\sum_{i=1}^{T-1}\left\|\Delta W_{T}^{\top}\Delta W_{i}\right\|_{2}$ (8)
(\@slowromancapv@) Generalized Sparsity Rate (GSR) Hurley and Rickard (2009): This metric quantifies the sparsity of a matrix when none of its elements are zero. For detailed information and proof, see Appendix 7.6. The formula is:
$S_{q,p}^{*}(\Delta W_{i})$ $=(mn)^{1/p-1/q}\frac{\ell_{q}(\Delta W_{i})}{\ell_{p}(\Delta W_{i})},(p>q>0)$ (9)
where $m\times n$ is the dimension of $\Delta W_{i}$, and $\ell_{p}$ and $\ell_{q}$ are the respective norms. In our case, $p=2$ and $q=1$.
(\@slowromancapvi@) Average Collision Rate (ACR): The concept of parameter collision was mentioned in  Kirkpatrick et al. (2017). Here, we formally define the evaluation metric ACR, which quantifies the Collision Rate (CR) between parameters across tasks. It is defined as:
$ACR_{T}=\frac{2}{T(T-1)}\sum_{1\leq i<j\leq T}CR_{i,j}$ (10)
where $CR_{i,j}$ is the collision rate between the $i$-th and $j$-th task’ parameters, which is the fraction of elements that are non-zero at the same positions in both matrices, as defined in Definition 1:
$CR_{i,j}=\frac{\sum_{a,b}\mathbf{1}\left(\Delta W_{i}[a,b]\neq 0\land\Delta W_{j}[a,b]\neq 0\right)}{n\times m}$ (11)
Table 1: Summary of results on two standard continual learning benchmarks using the T5-large model. The average accuracy after training on the final task is reported, with all results averaged over three independent runs.
Benchmarks Standard CL Benchmark Large Number of Tasks
Methods Order-1 Order-2 Order-3 avg Order-4 Order-5 Order-6 avg
ProgPrompt 75.2 75.0 75.1 75.1 78.0 77.7 77.9 77.9
PerTaskFT 70.0 70.0 70.0 70.0 78.1 78.1 78.1 78.1
MTL 80.0 80.0 80.0 80.0 76.5 76.5 76.5 76.5
SeqFT 18.9 24.9 41.7 28.5 7.4 7.4 7.5 7.4
SeqLoRA 44.6 32.7 53.7 43.7 0.6 1.9 1.6 1.6
IncLoRA 66.0 64.9 68.3 66.4 63.3 58.5 61.7 61.2
Replay 55.2 56.9 61.3 57.8 55.0 54.6 53.1 54.2
EWC 48.7 47.7 54.5 50.3 45.3 44.5 45.6 45.1
LwF 54.4 53.1 49.6 52.3 50.1 43.1 47.4 46.9
L2P 60.3 61.7 61.1 60.7 57.5 53.8 56.9 56.1
LFPT5 67.6 72.6 77.9 72.7 70.4 68.2 69.1 69.2
O-LoRA 75.4 75.7 76.3 75.8 72.3 64.8 71.6 69.6
N-LoRA $\bm{79.2}$ $\bm{78.4}$ $\bm{78.8}$ $\bm{78.8}$ $\bm{73.6}$ $\bm{70.3}$ $\bm{73.2}$ $\bm{72.4}$
Baselines. To demonstrate the effectiveness of N-LoRA, we follow the setup of O-LoRA  Wang et al. (2023a) and compare it against the following baseline methods:
(i) Independent Training Methods: PerTaskFT trains a separate model for each task, while ProgPrompt Razdaibiedina et al. (2023) learns independent prompts for each task, which essentially results in training a distinct model for each task.
(ii) Upper Bound: MTL represents multi-task learning, where a single model is trained on all tasks simultaneously. This method serves as the upper bound for continual learning.
(iii) Non-Continual Learning Methods: SeqFT de Masson D’Autume et al. (2019) fine-tunes all model parameters across a sequence of tasks. SeqLoRA trains a single LoRA adapter across the task sequence while freezing the pretrained model parameters. IncLoRA trains a new LoRA module for each task.
(iv) Continual Learning Methods: Replay fine-tunes all model parameters while replaying samples from previous tasks using a memory buffer. EWC Kirkpatrick et al. (2017) applies a regularization loss to adjust all model parameters. LwF Li and Hoiem (2017) preserves knowledge from previous tasks by recording their responses on new task data and using these responses as a regularization term. Both L2P Wang et al. (2022b) and LFPT5 Qin and Joty (2021) employ prompts, dynamically selecting or generating prompts to adapt to new tasks. O-LoRA Wang et al. (2023a) incrementally updates parameters for new tasks by constraining them to the orthogonal subspace of previous tasks while keeping the LoRA parameters learned from prior tasks fixed.
4.2 Comparisons with SOTA methods
Tables 1 and 2 present the performance comparison (Average Accuracy $A_{T}$) between N-LoRA and other baselines on different benchmarks. Following LFPT5 Qin and Joty (2021) and O-LoRA Wang et al. (2023a), we report the results from three independent runs using different task orders and models on the CL benchmark.
[…]
Table 2: Performance comparison between N-LoRA and O-LoRA on the larger LLaMA-7B model, reporting average accuracy across all task orders. Results are averaged over three independent runs.
Benchmarks Order-1 Order-2 Order-3 avg
O-LoRA 76.8 75.7 75.7 76.1
N-LoRA $\bm{77.2}$ $\bm{77.3}$ $\bm{78.4}$ $\bm{77.6}$
Performance on a Large Number of Tasks. Continual learning with a large number of tasks presents a significantly more challenging benchmark. As shown in Table 1, when the number of tasks increased, N-LoRA consistently outperforms all existing SOTA methods across all task orders (4, 5 and 6) in terms of average performance, with an average improvement of $\bm{2.8\%}$. However, despite this, N-LoRA, like other continual learning methods, still lags behind independent training approaches such as ProgPrompt and PerTaskFT, suggesting that there is considerable room for improvement in tackling a large number of tasks.
4.3 Enhancing Generalization of LLMs
This subsection explores how N-LoRA affects the generalization ability of LLMs, evaluated on the Standard CL Benchmark and Large Number of Tasks Benchmark. Using both N-LoRA and O-LoRA(current SOTA method), we conduct continual learning across the first $T-1$ tasks, followed by testing on the unseen $t$-th task. For comparison, we also evaluate pre-trained LLMs directly on the unseen task.
[…]
Figure 4: Performance comparison on unseen tasks on Average Accuracy metric. N-LoRA consistently outperforms O-LoRA across all task order settings. The pre-trained model performs poorly, with accuracy close to random on unseen tasks.
4.4 N-LoRA vs O-LoRA: Reducing Forgetting
In this subsection, we conduct an in-depth analysis of how N-LoRA reduces forgetting from multiple perspectives, providing a comprehensive comparison with O-LoRA, the current SOTA method.
Table 3: Performance comparison between O-LoRA and N-LoRA on Orthogonal Overlap (OO) and Adapter Weight Orthogonality Magnitude (AWOM) metrics. Lower values indicate better orthogonality.
Metrics OO()($\downarrow$)( ↓ ) AWOM()($\downarrow$)( ↓ )
Order O-LoRA N-LoRA O-LoRA N-LoRA
1 21.00 $\bm{3.80}$ 55.42 $\bm{0.13}$
2 20.50 $\bm{4.16}$ 126.25 $\bm{0.17}$
3 20.13 $\bm{3.42}$ 118.75 $\bm{0.15}$
4 30.75 $\bm{13.94}$ 14.40 $\bm{0.19}$
5 31.38 $\bm{11.06}$ 14.40 $\bm{0.19}$
6 34.50 $\bm{2.47}$ 6.57 $\bm{0.03}$
avg 26.38 $\bm{6.47}$ 55.96 $\bm{0.14}$
Higher Orthogonality. We adopt Orthogonal Overlap (OO) Wang et al. (2023a) and Adapter Weight Orthogonality Magnitude (AWOM) Bhardwaj et al. (2024) as metrics to assess orthogonality. As shown in Table 3, N-LoRA surpasses O-LoRA across both metrics, achieving ×$\bm{\times 4.07}$ times and ×$\bm{\times 388.3}$ times improvements, respectively. These results further demonstrate that the low collision rate induced by sparsity constraints can achieve stronger orthogonality compared to orthogonality constraints. Notably, although O-LoRA directly optimizes orthogonality using Orthogonal Overlap(OO) as its loss function, its actual OO is still lower than that of N-LoRA. This can likely be attributed to significant overlap in the task subspaces, leading to unavoidable conflicts, despite the orthogonal constraint. In contrast, as discussed in Section 3.3, N-LoRA reduces the dimensionality of task subspaces, decreasing the collison rate and consequently enhancing orthogonality.
Lower Sparsity Leads to Fewer Collisions and Reduced Forgetting. Figure 5 illustrates the relationship between sparsity, collision rate, and forgetting rate for N-LoRA and O-LoRA. It can be observed that the data points for N-LoRA are clustered in the lower-left corner of the plot, with an average forgetting rate of $\bm{2.37\%}$. In contrast, the data points for O-LoRA are in the upper-right corner, with an average forgetting rate of $\bm{4.20\%}$. This indicates that across varying benchmarks and orders, lower sparsity leads to fewer collisions and consequently reduces forgetting. The sparsity constraint in N-LoRA effectively minimizes collision rates, directly contributing to lower forgetting. Conversely, even though O-LoRA applies an orthogonality constraint on subspaces, it still encounters a high collision rate, resulting in greater forgetting.
Table 4: Performance comparison of O-LoRA and O-LoRA+N-LoRA on two benchmarks. O-LoRA+N-LoRA represents the integration of N-LoRA into O-LoRA, enhancing the performance of O-LoRA across all task orders.
Benchmarks Standard CL Benchmark Large Number of Tasks
Methods Order-1 Order-2 Order-3 avg Order-4 Order-5 Order-6 avg
O-LoRA 75.4 75.7 76.3 75.8 72.3 64.8 71.6 69.6
O-LoRA + N-LoRA $\bm{79.7}$ $\bm{79.8}$ $\bm{79.1}$ $\bm{79.5}$ $\bm{72.9}$ $\bm{67.2}$ $\bm{74.0}$ $\bm{71.4}$
Figure 5: The relationship between sparsity, collision rate, and forgetting. The x-axis and y-axis represent the Generalized Sparsity and Average Collision Rate, respectively. Data points for N-LoRA (red) are clustered in the lower-left corner, while O-LoRA (blue) data points are concentrated in the upper-right corner. The dashed ellipses indicate the average forgetting rates for each method.
[…]
Figure 6: Ablation study comparing different sparsity constraint methods. NS refers to no sparsity constraint, while SA and SB apply sparsity constraints only to LoRA matrices $A$ and $B$, respectively. SAB applies independent sparsity constraints to both matrices $A$ and $B$. Ours represents the N-LoRA method, which significantly outperforms all other methods.
4.5 Ablation Studies
In this subsection, we perform ablation studies to investigate the impact of sparsity constraints and the various ways of applying them. As illustrated in Figure 6, the model experiences severe forgetting when no sparsity constraint is applied (NS method). Applying sparsity to only LoRA matrix $A$ or matrix $B$ (SA and SB methods) can partially enhance parameter updates, resulting in some performance gains. However, since $\Delta W$ is the product of both $A$ and $B$, constraining only one matrix still permits considerable redundancy in the overall update. Simultaneously applying constraints to both $A$ and $B$ (SAB method) also fails to achieve optimal results. This is because the actual update during forward propagation is $\Delta W=AB$, and constraining the matrices independently hampers effective gradient flow, thereby impeding global optimization. The experimental results clearly show that only the N-LoRA approach (Ours) effectively reduces parameter collisions and substantially improves performance.
5 Conclusion
In this paper, we theoretically and empirically reveal that constructing non-collision parameters is a more critical factor in overcoming continual learning challenges. Building on this insight, we propose N-LoRA, a simple yet effective approach that leverages sparsity and low collision rates to improve continual learning. Extensive experiments on continual learning benchmarks demonstrate that, compared to SOTA methods, N-LoRA not only achieves better performance, higher task orthogonality, and lower parameter collision rates through sparsity, but also effectively reduces the dimensionality of task subspaces.
6 Limitations
Although N-LoRA demonstrates superior performance across multiple metrics in continual learning benchmarks compared to previous methods, several limitations remain. First, in scenarios involving a significantly larger number of tasks (e.g., hundreds or even thousands), the parameter space, despite the imposed sparsity constraint, can still become exhausted, leading to parameter collisions. The scalability and effectiveness of N-LoRA under such conditions require further investigation. Additionally, while N-LoRA performs well in controlled benchmark environments, its effectiveness in highly dynamic and more complex real-world continual learning tasks remains to be fully evaluated.